Christopher J. Fearnley

4807 Chester AVE APT 207 cjf@netaxs.com
Philadelphia PA 19143–3441 cjf@i21.com
(215)724–2265  http://www.netaxs.com/˜ cjf

Education

∘ Diagnostic Teaching (graduate) Drexel University, 1990. ∘ BA Mathematical Sciences/Philosophy SUNY–Binghamton, 1989. ∘ Regents Scholar, 1985–1989.

Computer Experience

∘ Chief of Engineering, Internet21. 1997–Present. ∘ Internet Engineering Consulting for several Philadelphia area Internet companies (including Net Access,

Nothin But Net, CrafTech Computer Solutions, Netlisters, OnIt Communications, CyberLoft Café, Cyber World Net, and Jaguar Netlink). Services provided include network engineering, web site development, web site CGI programming, system administration, Internet administration training, and technical support for end-users. 1995–Present.

∘ Member of the Debian GNU/Linux Project’s development team, 1995–Present. ∘ Founding member of the Philadelphia Linux User’s Group, 1996–Present. ∘ Author of a semi-regular column on UNIX in the Data Bus, 1993–1997. ∘ UNIX SIG leader and speaker, Philadelphia Area Computer Society, 1993–1996. ∘ Student Intern: System Design in LOTUS 1–2–3, NYS DOT Engineering R & D Bureau, Summer 1988.

Organizational Experience

∘ Promotion Analysis Administrator (customer and computer service), SmithKline Beecham, 1994. ∘ Bookkeeper for WPMF Food Co-op, 1993–Present. ∘ Support in preparation of FDA NDA for Tagamet-CR^{R}, SmithKline Beecham, 1992–1993. ∘ Retail Work: Supervisor, 1992; Prepare book displays, Cashier, 1989–1992. ∘ Organizer and lecturer for the Bucky Fuller Spaceship Earth Club, Binghamton, 1989. ∘ Voting Member Undergraduate Curriculum Committee, Math Department, Binghamton, 1986–1989. ∘ Member, V.P., President Math Club, Binghamton, 1985–1989. ∘ Organizer, President Chess Club, Binghamton, 1987–1989.

Teaching Experience

∘ Substitute Teacher, West Philadelphia Elementary Schools, 1990–1991. ∘ Volunteer Shop Teacher’s Aide (Elementary), 1989–1990. ∘ Chess Teacher, lessons to first graders, 1984-1985.

R. Buckminster Fuller Related Work

∘ Studies in Fuller’s life and work including Synergetics, Spherical Geometry, Critical Path Initiatives, Global

Democracy, Education, etc., I have a comprehensive interest in Fuller’s work, 1985–Present. ∘ Compiler and maintainer of the 105 page “Bucky Fuller FAQ” distributed on the Internet, 1994–Present. ∘ Major contributor to FIX (Fuller Information eXchange) BBS message base, 1991–1994. ∘ Active participant on GEODESIC, the list for the discussion of R. Buckminster Fuller. 1989–Present. ∘ Published article: “Reading Synergetics: Some Tips” in Trimtab, Autumn 1991.